# LLM-Guided Open RAN: Empowering Hierarchical RAN Intelligent Control

### v1

@@ IV Future Challenges
In this section, we discusses the future challenges of LLMs in O-RAN. As shown in Fig. 4, key challenges include leveraging multi-modal data, fine-tuning of LLMs, enabling systematic collaboration between RICs, and achieving timely inference. Each topic is discussed in detailed below.
Figure 4: Future challenges.
IV-A Leveraging Multi-Model Data
Multi-modal data (e.g. text, signals, images, and videos) can be used for better understanding on the complex network environments. The multi-modal LLMs are capable of reasoning across various data types, providing richer contextual insights and enabling more informed decision-making.
Meanwhile, the zero-shot or few shot learning methods (e.g. retrieval augmentd generation (RAG) and prompt engineering) have been explored to adapt LLMs to new environments. However, several open questions remain. First, current prompt engineering methods generally lack support for multi-modal inputs in wireless communication domain, making it critical to develop methods that effectively handle such data. Moreover, the performance of RAG is highly depend on retrieval engine. Designing an effective multi-modal RAG engine in wireless communication domain is key challenge for improving LLM reasoning.
@@ IV-B Fine-Tuning of LLMs
While current LLMs have been trained well for general domain, they are not specifically optimized for wireless communication tasks, which limits their performance in this area. Although, prompt engineering can guide LLMs to some extent, limitations in context window size and the dynamic nature of wireless environment hinder real time reasoning.
To address this issue, fine-tuning LLMs using offline, domain-specific data is essential. Fine-tuning LLMs can enhance decision-making relevance and reliability. Nonetheless, two major open challenges remain: 1) how to collect or effectively utilize limited wireless communication data, and 2) how to effectively generate high quality instruction data for fine-tuning?
@@ IV-C Systematic Collaboration between RICs
Due to the lack of domain specific fine-tuning, LLMs-generated policies are not always optimal. Enhancing them with lightweight AI techniques is critical, especially for near-RT applications where inference speed is paramount. To fully exploit the strength of both LLMs (for long time, global optimization) and lightweight AI (for real-time responsiveness), a cooperative framework must be established. This framework should enable smooth interaction between the LLMs and lightweight AI models, supporting adaptive and robust decision-making under dynamic conditions.
Several challenges need to be addressed. First, LLMs operate on large time scale, while network environment change rapidly in near-RT. Bridging this time gap to make training and prediction work well is not easy. Second, due to the limited capacity of lightweight models, selecting the most relevant features for specific tasks is critical for performance enhancement.
@@ IV-D Timely Inference
Modern LLMs, such as GPT-4 with 1.7 trillion parameters, demand significant computational resources and are often deployed on cloud-based servers. This dependency introduces substantial latency, making deploying LLM model in near-RT RIC as a xApp challenging.
To address this challenge, technologies such as model pruning, quantization and knowledge distillation are crucial for reduce the model size. Specifically, the quantization technology reduces the precision of model parameters to lower computational overhead without significantly affecting performance. In addition, the model pruning and knowledge distillation train a smaller, optimized versions of LLMs. The different methods can compress the model with different ratio. However Compression methods typically trade off some model performance for reduced latency. Balancing model size, inference speed, and accuracy is a critical challenge.
@@ V Conclusions
This paper proposes the LLM-hRIC framework for O-RAN. In this framework, the LLM is deployed as rApp in the non-RT RIC to provide strategic guidance, while RL is implemented as xApp in the near-RT RIC to facilitate near-RT decision-making. The superior performance highlight the effectiveness of combining LLM-empowered non-RT RIC and RL-empowered near-RT RIC to tackle the dynamic and complex resource management. Meanwhile, this work highlights the potential of integrating global reasoning and local responsiveness, establishing the LLM-hRIC framework as a foundation for scalable, adaptive, and high-performance solutions in future O-RAN systems.

### v2

@@ IV Future Challenges
In this section, we discusses the future challenges of LLM-hRIC for O-RAN. As shown in Fig. 5, key challenges specific to LLM-hRIC include leveraging multi-modal data, fine-tuning LLMs for guidance generation, enabling systematic collaboration between RICs in LLM-hRIC, and achieving timely inference. Each topic is discussed in detailed below.
Figure 5: Future challenges of the LLM-hRIC for O-RAN.
IV-A Leveraging Multi-Model Data in LLM-hRIC
In the LLM-hRIC, the multi-modal LLMs can reason across various data types. Multi-modal data (e.g., text, signals, images, and videos) provides detailed environment insights, enhancing the ability of the non-RT RIC to generate more accurate and context-aware guidance for near-RT RIC. Moreover, zero-shot and few-shot learning techniques, such as prompt engineering and retrieval augmentd generation (RAG), have been explored to enhance the performance of LLMs in new environments. Despite their potential, significant challenges remain: 1) prompt engineering lacks robust support for multi-modal inputs in the wireless communication domain, necessitating new methods to handle such data for generating high-quality guidance. 2) RAG performance relies heavily on the retrieval engine. Designing an effective multi-modal RAG engine for LLM-hRIC in the wireless communication domain is critical to improve LLM reasoning and guidance generation.
@@ IV-B Fine-Tuning of LLMs
Current LLMs are trained for general domains but not optimized for wireless communication tasks, limiting guidance quality in LLM-hRIC. While prompt engineering can partially bridge this gap, its effectiveness is limited by the quality of examples and input token window size. Fine-tuning LLMs with offline, domain-specific data is essential to enhance guidance reliability in the non-RT RIC. However, two challenges remain: 1) Collecting or effectively utilizing the limited available data on wireless communication is challenging. 2) Generating high-quality instruction data for fine-tuning is essential. Automated or semi-automated methods for generating diverse, domain-relevant instruction data must be developed to improve the guidance quality of fine-tuned LLMs.
@@ IV-C Systematic Collaboration between RICs
In LLM-hRIC, the LLMs-empowered non-RT RIC generates guidance for multiple near-RT RICs, each responsible for diverse tasks. A systematic collaboration framework is needed to ensure smooth interaction between the LLMs and lightweight AI models. Several challenges need to be addressed. 1) non-RT RICs operates on a large time scale, while near-RT RICs can operate on different small time scales. This temporal mismatch complicates the synchronization between the guidance generation in the non-RT RIC and the real-time decision-making in the near-RT RIC. 2) near-RT RICs perform a variety of tasks (e.g., power allocation, PRB allocation, and user association). Each task requires specific guidance tailored to its unique requirements, making it crucial to develop a mechanism for providing task-aware, customized guidance for different near-RT RICs.
@@ IV-D Timely Inference
Modern LLMs, such as GPT-4 with 1.7 trillion parameters, require significant computational resources and are often cloud-deployed, causing latency issues. This makes deploying an LLM in the non-RT RIC challenging, as some non-RT RIC require latency near 1s. Technologies like model pruning, quantization, and knowledge distillation can reduce model size. Specifically, the quantization technology reduces the precision of model parameters to lower computational overhead, while pruning and distillation create smaller, optimized LLMs. These methods compress models at varying ratios but often trade off performance for speed. Balancing model size, inference speed, and accuracy is critical to ensuring efficient deployment of LLMs in the LLM-hRIC framework.
@@ V Conclusions
This paper proposes the LLM-hRIC framework for O-RAN. In this framework, the LLM-empowered non-RT RIC provides the strategic guidance, while the RL-empowered near-RT RIC explores guidance with local observation to facilitate near-RT decision-making. The superior performance highlight the effectiveness of combining LLM-empowered non-RT RIC and RL-empowered near-RT RIC to tackle the dynamic and complex resource management. Meanwhile, this work highlights the potential of integrating global reasoning and local responsiveness, establishing the LLM-hRIC framework as a foundation for scalable, adaptive, and high-performance solutions in future O-RAN systems. Additionally, we identify key challenges for LLM-hRIC, as discussed in Section IV, to guide future research and development.